Determine the net heat transfer through the black body disc with radius 4cm and mass 9g, at 320C when transferred to a surrounding temperature of 650C.
1
Expert's answer
2020-12-15T11:46:23-0500

Answer

the net heat transfer through the black body disc is given by

"\frac{dQ_{in}}{dt}=mc\frac{dT}{dt}"

So heat transffer per unit second

"dQ=mc\space dT"

Using given data

"dQ=0.009\times0.385\times(330+273) \\=2.09J"